# Causality Pursuit from Heterogeneous Environments via Neural Adversarial Invariance Learning

### v1

@@ 6.2.2 Finite Performance of FAIR-NN Estimator
Figure 4: The visualization of (a) the SCM and (b) the ${\mathrm{sig}(w)}$ during training in one trail for FAIR-NN estimator when $k=1$. We use different colors to represent the different relationships with $Y$: blue = parent, red = child, orange = offspring, lightblue = other.
@@
as the evaluation metric. We report the median of $\widehat{\mathtt{MSE}}$ over $100$ replications for the estimators (1) – (4) akin to that for the linear model. For (1), (2), and (4), we also use ReLU neural network width depth $2$ and width $128$ in running least squares. Fig. 4 (b) also visualizes how the Gumbel gate values for different covariables ${\mathrm{sig}(w)}$ evolve during training in one trail. We can see that the training dynamics for ${\mathrm{sig}(w)}$ is much more challenging and interesting than that for the linear model depicted in Fig. 2: the weight for some $Y$’s children quickly increases at a comparable rate than the variables in $S^{\star}$ at the beginning, but such a trend slows down and finally completely reverses in the middle. We leave the rigorous and in-depth analysis behind such dynamics for future studies.
@@
    contains the omitted discussions in the main text, including the applicable scenarios for the nonparametric invariance pursuit, some discussions and extensions on the method, and some discussions on the conditions in Section 4.
@@
    contains the complete result that is sketched in Section 5.
@@ A.1 Applicable Scenarios for Nonparametric Invariance Pursuit
The SCM (4.4) and Proposition A.1 extend the framework described in Peters et al., (2016) (specifically Section 4.1 and Proposition 1). This model accommodates nonlinear structural assignments. Critically, the residuals $\varepsilon^{(e)}=Y^{(e)}-\mathbb{E}[Y^{(e)}|X_{S^{\star}}^{(e)}]$, do not need to be independent of $X_{S^{\star}}^{(e)}$ or remain invariant across various environments as represented by $\varepsilon^{(e)}\sim\mu_{\varepsilon}$. Such flexibility broadens the scope for various applications, including binary classification. According to Proposition A.1, when restricted to model (4.4), a specific instantiation of our generic statistical model (1.1), identifying the true important variable set $S^{\star}$ is tantamount to pinpointing the direct cause of the target variable $Y$. Concurrently, unveiling the invariant association $m^{\star}$ aligns with uncovering the causal mechanism between $Y$ and its direct causes.
@@ A.2 Discussion on the Methods
[Q] You are doing “focused regularizer” that are of combinatorial nature, can it be removed?
@@ A.3 Extensions to General Environment Variable and Loss Function
In the main text, we propose an estimation framework leveraging conditional expectation invariance with respect to discrete environment variables. It is worth noticing that our adversarial estimation framework is indeed more versatile than this: one can easily extend it to other conditional point prediction invariance with respect to more general environment covariate. We briefly discuss the direct extension here and leave a rigorous treatment as future work. In the following discussions, suppose we observe data $\{(X_{i},Y_{i},E_{i})\}_{i=1}^{n}$ drawn i.i.d. from some distribution $\mu_{0}$, where $X\in\mathbb{R}^{d}$ is the covariate we used for prediction, $Y\in\mathbb{R}$ is the target response, $E\in\mathbb{R}^{q}$ is the environment covariate we wish our prediction should be invariant with respect to.
@@
where $\gamma$ is the hyper-parameter to be determined, and $\mathcal{F}_{S_{g}}=\{f(x,e)\in\mathcal{F},f(x,e)=w(x_{S_{g}},e)\text{ for some }w\}$. Similar to the calculation for (3.3), one can expect that minimizing the population counterpart of the focused adversarial invariance regularizer $\max_{f\in\mathcal{F}_{S_{g}}}\widehat{\mathsf{J}}(g,f)$ shares a similar nature of imposing (A.1). One can derive non-asymptotic identification and estimation error results akin to Theorem B.2 and Theorem B.3 provided strong convexity and certain Lipschitz property of the loss $\ell(u,y)$. We leave this for future studies.
@@ A.4 Discussion on Relaxing Nonparametric Invariance Pursuit Identification Condition
The above condition is definitely weaker than 4.5 because 4.5 essentially requires the set $S^{\star}$ is the maximum invariant set,
@@
Here (A.4) just rewrites 4.5 in a manner similar to (A.3).
It is easy to derive results similar to Theorem 4.1 under (A.3) rather than 4.5. We can construct cases where (A.3) holds but 4.5 does not. Examples include Example 1 below with $s^{(1)}s^{(2)}=1$ under which both $\{1\}$ and $\{2\}$ are invariant set but the set $\{1,2\}$ is not. In this case, 4.5 no longer holds. However, our algorithm can still consistently estimate $m^{\star}$ provided (A.3) holds, that the variable $X_{1}$ has better prediction power. In the main text, we still adopt 4.5 instead of (A.3). The main reasons are as follows. All our discussions are under the SCM with interventions setting.
Firstly, as further shown in Section 4.2, the cases where 4.5 fails to hold are degenerate cases. When the interventions are nondegenerate, there always exists a maximum invariant set $S^{\star}$, i.e., 4.5 holds. This means, (A.3) is somewhat “marginally” weaker than 4.5.
The second reason is the lack of semantic meaning of $S^{\star}$ under this case. When the interventions are non-degenerate, $S^{\star}$ can be interpreted as “contemporary/pragmatic direct causes” that can be expressed as direct causes + unaffected children + parents of unaffected children in Proposition A.2, such a variable set also has certain robust transfer learning properties as stated in Proposition 4.4. All the above semantic meanings are valid even if the interventions are insufficient. However, when the interventions are degenerate such that 4.5 may not hold but (A.3) may hold, e.g., Example 1, all the two properties will no longer hold. If the true causal mechanism is $X_{1}\to Y\to X_{2}$, then it is possible to construct data generating process such that $S^{\star}$ can be either $\{1\}$ or $\{2\}$ in (A.3).
@@ A.5 Discussion on the Nondegenerate Intervention Condition
The conditions (1) and (2) in 4.8 are imposed to eliminate some degenerate cases. To illustrate the intuitions why such two conditions are needed, and how such a condition will hold in general. We consider the following two examples.
Introduction of condition (1)
From a high-level viewpoint, the introduction of condition (1) is to eliminate the cases where though there are shifts in condition distributions among different environments, it happens that there are no shifts in conditional expectations. This can be illustrated in the following example.
@@
where $U_{1},U_{2},U_{3}$ are independent standard normal variables, and =$\mathcal{E}=\{1,2\}$. We let $e=1$ be the observational environment and $e=2$ be the intervention environment where the linear effect of $Y$ on $X_{2}$ are intervened ($s^{(1)}\neq s^{(2)}$). We also focus on the regime where $s^{(1)}+s^{(2)}\neq 0$ such that running least squares will lead to a biased solution.
@@
The introduction of 4.8 (1) is to rule out the cases where $s^{(1)}=1/s^{(2)}=s$. And it is easy to see when $s^{(1)}$ and $s^{(2)}$ are independently generated from some prior distribution that is absolute continuous with respect to Lebesgue measure on $\mathbb{R}$, i.e., $S^{(1)},S^{(2)}\sim p_{s}$, then
@@
Introduction of condition (2).
The condition (2), that the faithfulness condition on $\widetilde{M}$, is to eliminate the cases where though the interventions are applied, it happens that such interventions do not make an impact on the variables intervened. The following example presents such an example.
@@
where $U_{1},U_{2},U_{3}$ are independent standard normal variables, $s\neq 0.5$ is a fixed parameter. We let $e=1$ be the observational environment and $e=2$ be the intervention environment where shifts in mean are applied to the variables $X_{1}$ and $X_{2}$.
@@
A.6 The Complete Statement of Proposition 4.4

### v2

@@ 5.2.2 Finite Performance of FAIR-NN Estimator
Figure 4: The visualization of (a) the SCM and (b) the ${\mathrm{sig}(w)}$ during training in one trial for FAIR-NN estimator when $k=1$. We use different colors to represent the different relationships with $Y$: blue = parent, red = child, orange = offspring, lightblue = other.
@@
as the evaluation metric. We report the median of $\widehat{\mathtt{MSE}}$ over $100$ replications for the estimators (1) – (4) akin to that for the linear model. For (1), (2), and (4), we also use ReLU neural network width depth $2$ and width $128$ in running least squares. Fig. 4 (b) also visualizes how the Gumbel gate values for different covariables ${\mathrm{sig}(w)}$ evolve during training in one trial. We can see that the training dynamics for ${\mathrm{sig}(w)}$ is much more challenging and interesting than that for the linear model depicted in Fig. 2: the weight for some $Y$’s children quickly increases at a comparable rate than the variables in $S^{\star}$ at the beginning, but such a trend slows down and finally completely reverses in the middle. We leave the rigorous and in-depth analysis behind such dynamics for future studies.
@@
5.3 Application I: Discovery in Real Physical Systems
We apply our method to perform causal discovery in the light tunnel datasets from Gamella et al., (2024). The data are collected from a real physical device under different manipulation settings. The tunnel device contains a controllable light source at one end and two linear polarizers mounted on rotating frames. Several sensors are deployed in various positions to measure the light intensity. The causal relationships between the variables of interest is known such that we can get access to the ground-truth cause-effect relationship; see Fig. 2(d) and Fig. 3(a) therein for the device diagram and the cause-effect graphs, respectively. It is worth noticing that the data are collected from a real-world device where the associations between the measurements follow from real-world physical laws. This realistic nature together with the knowledge of ground-truth cause-effect knowledge make it an excellent testbed for causal discovery algorithms.
Using the notation in Gamella et al., (2024), we use the variables $(R,G,B,\theta_{1},\theta_{2},\widetilde{V}_{3},\widetilde{V}_{2},\widetilde{V}_{1},\widetilde{I}_{3},\widetilde{I}_{2},\widetilde{I}_{1},\widetilde{C})$. Here $(R,G,B)$ is the intensity of the light source at three different wavelengths, $\widetilde{C}$ is the drawn electric current, $(\theta_{1},\theta_{2})$ represent the angles of the polarizer frame, and $(\widetilde{V}_{3},\widetilde{V}_{2},\widetilde{V}_{1},\widetilde{I}_{3},\widetilde{I}_{2},\widetilde{I}_{1})$ are the measurement of light-intensity sensors in various positions.
We plan to learn algorithmically the direct cause for $Y=\widetilde{I}_{3}$, the infrared measurement of the light-intensity sensor after the polarizers, among a subset of manipulable variables and measurement variables $(X_{1},\ldots,X_{11})=(R,G,B,\theta_{1},\theta_{2},\widetilde{V}_{3},\widetilde{V}_{2},\widetilde{V}_{1},\widetilde{I}_{2},\widetilde{I}_{1},\widetilde{C})$ under the following two-environment experimental setting: $e=0$ is the observational environment, $e=1$ is the interventional environment where the variables $\{\widetilde{V}_{j}\}_{j=1}^{3}$ and $\{\widetilde{I}_{j}\}_{j=1}^{2}$ are weakly intervened on. This leads to the following “equivalent” ground-truth cause-effect relationship among those variables and the effect of “environment intervention” in Fig. 6 (a). In this case, the variables $(R,G,B,\theta_{1},\theta_{2})$ are the direct causes, i.e., $S^{\star}=\{1,2,3,4,5\}$, $\widetilde{V}_{3}$ are the spurious variables that will lead to biased estimation. The remaining variables are exogenous but have marginal predictive power, i.e., $\mathrm{Var}[Y|X_{j}]>0$ for $j\geq 7$.
We will use the following dataset in the experiment: the environment dataset $\mathcal{D}_{0}$ with size $|\mathcal{D}_{0}|=10^{4}$, the weakly interventional environment dataset $\mathcal{D}_{1}$ with $|\mathcal{D}_{1}|=3000$, and five strongly interventional environment dataset $\mathcal{D}_{2,Z}$ with $Z\in\{\widetilde{V}_{1},\widetilde{V}_{1},\widetilde{V}_{3},\widetilde{I}_{1},\widetilde{I}_{2}\}$ and $|\mathcal{D}_{2,Z}|=1000$. In each trial, different methods use the same random subsample $\breve{\mathcal{D}}=\{\breve{\mathcal{D}}_{0},\breve{\mathcal{D}}_{1}\}$ with $\breve{\mathcal{D}}_{k}\subseteq\mathcal{D}_{k}$ and $|\breve{\mathcal{D}}_{k}|=n=1000$ to fit the model. How the fitted model $\widehat{f}$ quantitatively depends on exogenous/endogeneous spurious variable $Z$ is evaluated using the OOS $R^{2}$ in corresponding $\mathcal{D}_{2,Z}$ defined as
$R^{2}_{\mathrm{OOS},Z}:=\frac{\sum_{(X,Y)\in\mathcal{D}_{2,Z}}\{\widehat{f}(X)-Y\}^{2}}{\sum_{(X,Y)\in\mathcal{D}_{2,Z}}\{Y-\bar{Y}\}^{2}}\qquad\text{with}\qquad\bar{Y}=\frac{\sum_{(X,Y)\in\breve{\mathcal{D}}_{0}\cup\breve{\mathcal{D}}_{1}}Y}{2n}.$
See the detailed data collection and experimental configuration in Section C.3.
[…]
Figure 6: Discovery in Real Physical Systems: (a) the unified cause-effect relationship and interventions similar to Fig. 1 (b). (b) the average out-of-sample $R^{2}$ for different estimators using the spider chart: the axis annotated by placeholder variable $Z$ corresponds to the test environment where $Z$ is strongly intervened on. We can see the performance of Oracle-NN and FAIR-NN-RF is almost identical. (c) the average (based on 100 replications) of the worst-case (across 5 environments) of OOS $R^{2}$ for different methods as a function of $n$. (d) the variable selection rate over $100$ trials for different methods (top panel) and the variable selection rate for FAIR-NN for various $n$ (bottom panel). We use different colors to represent different relationships with $Y$: blue=parent, red=child, orange=neither ancestor nor descendants. (e) the distribution of worst-case OOS $R^{2}$ (y-axis) for Gumbel-trick optimized FAIR-NN (Gumbel), the follow-up refitted estimator (Refit), and Pooled LS (Pooled) when FAIR-NN selects the wrong variables: the subplots from top to bottom consider the cases of (i) failure in selection consistency (ii) false positive that it falsely selects the child $X_{8}=\widetilde{V}_{3}$ (iii) false negative that it does not select the entire ground-truth $(X_{1},\ldots,X_{5})=(R,G,B,\theta_{1},\theta_{2})$.
The first four rows in Fig. 6 (d) report the variable selection result for several methods over $100$ trials. The nonlinear ICP (Heinze-Deml et al.,, 2018) method does not select any variables because of its conservative nature and stronger heterogeneity condition to recover the direct cause. We can see that FAIR-NN can successfully recover the direct cause $(R,G,B,\theta_{1},\theta_{2})$ in this case. It exploits neural networks’ capability in efficiently detecting the nonlinear associations (the Malus’s law, $\widetilde{I}_{3}\propto\cos^{2}(\theta_{1}-\theta_{2})$ for fixed $(R,G,B)$), while the linear counterpart FAIR-Linear fails to select the variables $(\theta_{1},\theta_{2})$. It is worth pointing out that such a causality recovery cannot be attained by the traditional predictive power and simplicity tradeoff: the variable selection method based on random forest variable importance measures (ForestVarSel) in Heinze-Deml et al., (2018) cannot detect $(G,B,\theta_{1},\theta_{2})$ and falsely select $(\widetilde{I}_{1},\widetilde{I}_{2})$. The last three rows in Fig. 6 (d) illustrate how the variable selection rate for the FAIR-NN estimator changes when $n$ grows.
Fig. 6 (b) offers a quantitative illustration by showing the out-of-sample (OOS) $R^{2}$ of different estimators under environments with strong interventions on $(\widetilde{I}_{1},\widetilde{I}_{2},\widetilde{V}_{1},\widetilde{V}_{2},\widetilde{V}_{3})$, respectively. The estimator denoted as Oracle-$M$ with $M\in\{\mathrm{Linear},\mathrm{NN}\}$ referred to the method that runs regress $Y$ on $X_{S^{\star}}$ using model $M$. In the spider chart, the red shade represents the out-of-sample $R^{2}$ under different interventions for the Oracle-NN estimator that regresses $Y$ on its direct causes. We can see that its performances behave uniformly under various interventions: all the OOS $R^{2}$ are approximately equal to $0.91$. This is slightly better than that for the linear model ( Oracle-Linear) by 0.04. This illustrates the capability of neural networks introduced to detect weak, nonlinear causal signals from heterogeneous environments. The PoolLS-NN estimator regressing $Y$ on $X$ using neural network and all the data fully exploits the strong spurious association between $\widetilde{V}_{3}$ and $Y=\widetilde{I}_{3}$, its heavy reliance on $\widetilde{V}_{3}$ let it predict better (than the causal model Oracle-NN) when $\widetilde{V}_{3}$ is not intervened. However, its OOS $R^{2}$ significantly decreases by $0.2$ when $\widetilde{V}_{3}$ is strongly intervened hence the spurious association changes. On the contrary, the OOS $R^{2}$ for FAIR-NN after refitting ( FAIR-NN-RF) behaves almost identical to that for Oracle-NN. This quantitative result illustrates its capability to correct non-trivial and strong bias without no supervision and its efficiency in detecting nonlinear and weak signals.
Fig. 6 (c) shows how the worst-case OOS $R^{2}$ among the five, strong intervention environments changes for different estimators when $n$ grows. The performance of the Gumbel-trick optimized FAIR-NN estimator without refitting ( FAIR-NN-GB) lies between Oracle-NN and Oracle-Linear and significantly outperforms that of the PoolLS-NN estimator. This suggests that the gradient descent optimized algorithm has already found predictions nearly independent of the spurious variable, and the success of variable selection in Fig. 6 (d) is not because of truncating weak but non-negligible spurious signals. Moreover, as shown in Fig. 6 (e), its performance significantly outperforms the least squares estimator using either the full covariate or the selected covariates when it selects the wrong variable. This further supports the theoretical claims and the advantages of adopting penalized least squares.
5.4 Application II: Prediction Based on Extracted Features
We consider an image object classification task with a spurious background. The target is to classify water birds ($Y=1$) and land birds ($Y=0$) (see examples in Fig. 7 (a)) under backgrounds of water or land based on the feature $X\in\mathbb{R}^{500}$ extracted from ResNet pre-trained on ImageNet. We train a linear classifier on top of $X$ using data from two environments $\{\mathcal{D}_{k}\}_{k=1}^{2}$. In the first environment $\mathcal{D}_{1}$, $r_{w}=95\%$ water birds appear on the water background and $r_{l}=90\%$ land birds stay in land background. The spurious correlation numbers are $r_{w}=75\%$ and $r_{l}=70\%$ in $\mathcal{D}_{2}$. A good predictor should based on the core features related to the bird’s appearance rather than the strong spurious correlation between the background and label. The trained model is evaluated in a test environment $|\mathcal{D}_{3}|$ where the spurious correlation reverses: $r_{w}=2\%$ and $r_{l}=2\%$. We repeat the experiment $10$ times, where in each trial the training dataset and the test dataset are sampled from a larger dataset with sizes $n=|\mathcal{D}_{1}|=|\mathcal{D}_{2}|=50k$ and $|\mathcal{D}_{3}|=30k$. We compare our FAIR(-egulerized) estimator using $\mathcal{G}=\{{\mathrm{sig}(\beta^{\top}x)}\},\mathcal{F}=\{\beta^{\top}x\}$ and classification loss $\ell(y,v)=-\log(1-v)-y\log\{v/(1-v)\}$ ( FAIR-GB) with invariant risk minimization ( IRM) (Arjovsky et al.,, 2019) and group distributionally robust optimization ( GroupDRO) (Sagawa et al.,, 2020). We also consider running Lasso on different environments for reference, including (1) using all the data $\mathcal{D}_{1}\cup\mathcal{D}_{2}$ ( Pooled Lasso); (2) using data in $\mathcal{D}_{2}$ ( Lasso on D2); (3) using another randomized controlled environment $\mathcal{D}_{4}$ with $r_{w}=r_{l}=50\%$ and $|\mathcal{D}_{4}|=n$ ( Oracle). All the models are linear, and the performance of (3) can be seen as the upper bound of the performance using linear models; see data collection and experiential configuration details in Section C.4.
The performances are reported in Fig. 7 (b). Fig. 7 (c) also depicts how test accuracy changes as iterations in one trial. FAIR-GB performs similar to Oracle and significantly outperforms Lasso on D2, while other methods ( IRM, DRO) falls behind Lasso on D2. This indicates that these methods cannot go beyond interpolating the spurious associations in $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$, while our method can nearly eliminate the spurious association using the relatively small perturbations in the two environments.
MethodTest AccuracyOracle91.06 $\pm$ 0.24 %Lasso on D284.57 $\pm$ 0.71 %Pooled Lasso79.08 $\pm$ 0.54 %IRM80.32 $\pm$ 0.67 %GroupDRO82.75 $\pm$ 1.10 %FAIR-GB89.56 $\pm$ 0.53 %
Figure 7: Prediction Based on Extracted Features: (a) provides two sample images in the dataset: land bird on land (up) and water bird in water (bottom). (b) reports the average $\pm$ standard deviation of test accuracy over 10 trials for different estimators. (c) shows how the test accuracy changes over iterations for different methods in one trial.
Acknowledgement
We thank Yiran Jia for helpful discussions on presenting a generic identification result on SCM using the unified graph including $E$, Yimu Zhang for the help with the numerical implementation in Section 5.4, and Xinwei Shen for suggestions of using Gumbel approximation in implementation.
@@ Acknowledgement
    contains the omitted discussions in the main text, including the applicable scenarios for the nonparametric invariance pursuit, some discussions and extensions on the method, and some discussions on the conditions in Section 2 and Section 3.
@@
    contains the complete result that is sketched in Section 4.3.
@@ A.1 Applicable Scenarios for Nonparametric Invariance Pursuit
The SCM (3.2) and Proposition 3 extend the framework described in Peters et al., (2016) (specifically Section 4.1 and Proposition 1). This model accommodates nonlinear structural assignments. Critically, the residuals $\varepsilon^{(e)}=Y^{(e)}-\mathbb{E}[Y^{(e)}|X_{S^{\star}}^{(e)}]$, do not need to be independent of $X_{S^{\star}}^{(e)}$ or remain invariant across various environments as represented by $\varepsilon^{(e)}\sim\mu_{\varepsilon}$. Such flexibility broadens the scope for various applications, including binary classification. According to Proposition 3, when restricted to model (3.2), a specific instantiation of our generic statistical model (1.1), identifying the true important variable set $S^{\star}$ is tantamount to pinpointing the direct cause of the target variable $Y$. Concurrently, unveiling the invariant association $m^{\star}$ aligns with uncovering the causal mechanism between $Y$ and its direct causes.
@@ A.2 Discussion on the Methods
[Q] You are doing “focused regularizer” that are of combinatorial nature in computation, can it be removed?
@@ A.3 Extensions to General Environment Variable and Loss Function
In the main text, we propose an estimation framework leveraging conditional expectation invariance with respect to discrete environment variables. It is worth noticing that our adversarial estimation framework is indeed more versatile than this: one can easily extend it to other conditional point prediction invariance with respect to more general environment covariates. We briefly discuss the direct extension here and leave a rigorous treatment as future work. In the following discussions, suppose we observe data $\{(X_{i},Y_{i},E_{i})\}_{i=1}^{n}$ drawn i.i.d. from some distribution $\mu_{0}$, where $X\in\mathbb{R}^{d}$ is the covariate we used for prediction, $Y\in\mathbb{R}$ is the target response, $E\in\mathbb{R}^{q}$ is the environment covariate we wish our prediction should be invariant with respect to.
@@
where $\gamma$ is the hyper-parameter to be determined, and $\mathcal{F}_{S_{g}}=\{f(x,e)\in\mathcal{F},f(x,e)=w(x_{S_{g}},e)\text{ for some }w\}$. Similar to the calculation in Section 1.2, one can expect that minimizing the population counterpart of the focused adversarial invariance regularizer $\max_{f\in\mathcal{F}_{S_{g}}}\widehat{\mathsf{J}}(g,f)$ shares a similar nature of imposing (A.1). One can derive non-asymptotic identification and estimation error results akin to Theorem 4 and Theorem 5 provided strong convexity and certain Lipschitz property of the loss $\ell(u,y)$. We leave this for future studies.
@@ A.4 Discussion on Relaxing Nonparametric Invariance Pursuit Identification Condition
The above condition is definitely weaker than 2 because 2 essentially requires the set $S^{\star}$ is the maximum invariant set,
@@
Here (A.4) just rewrites 2 in a manner similar to (A.3).
It is easy to derive results similar to Theorem 1 under (A.3) rather than 2. We can construct cases where (A.3) holds but 2 does not. Examples include Example 1 below with $s^{(1)}s^{(2)}=1$ under which both $\{1\}$ and $\{2\}$ are invariant set but the set $\{1,2\}$ is not. In this case, 2 no longer holds. However, our algorithm can still consistently estimate $m^{\star}$ provided (A.3) holds, that the variable $X_{1}$ has better prediction power. In the main text, we still adopt 2 instead of (A.3). The main reasons are as follows. All our discussions are under the SCM with interventions setting.
Firstly, as further shown in Section 3, the cases where 2 fails to hold are degenerate cases. When the interventions are nondegenerate, there always exists a maximum invariant set $S^{\star}$, i.e., 2 holds. This means, (A.3) is somewhat “marginally” weaker than 2.
The second reason is the lack of semantic meaning of $S^{\star}$ under this case. When the interventions are non-degenerate, $S^{\star}$ can be interpreted as “contemporary/pragmatic direct causes” that can be expressed as direct causes + unaffected children + parents of unaffected children in Proposition 4, such a variable set also has certain robust transfer learning properties as stated in Proposition 2. All the above semantic meanings are valid even if the interventions are insufficient. However, when the interventions are degenerate such that 2 may not hold but (A.3) may hold, e.g., Example 1, all the two properties will no longer hold. If the true causal mechanism is $X_{1}\to Y\to X_{2}$, then it is possible to construct data generating process such that $S^{\star}$ can be either $\{1\}$ or $\{2\}$ in (A.3).
@@ A.5 Discussion on the Nondegenerate Intervention Condition
The conditions (a) and (b) in 5 are imposed to eliminate some degenerate cases. To illustrate the intuitions why such two conditions are needed, and how such a condition will hold in general. We consider the following two examples.
Introduction of condition (a)
From a high-level viewpoint, the introduction of condition (a) is to eliminate the cases where though there are shifts in condition distributions among different environments, it happens that there are no shifts in conditional expectations. This can be illustrated in the following example.
@@
where $U_{1},U_{2},U_{3}$ are independent standard normal variables, and =$\mathcal{E}=\{1,2\}$. We let $e=1$ be the observational environment and $e=2$ be the interventional environment where the linear effect of $Y$ on $X_{2}$ are intervened ($s^{(1)}\neq s^{(2)}$). We also focus on the regime where $s^{(1)}+s^{(2)}\neq 0$ such that running least squares will lead to a biased solution.
@@
The introduction of 5 (a) is to rule out the cases where $s^{(1)}=1/s^{(2)}=s$. And it is easy to see when $s^{(1)}$ and $s^{(2)}$ are independently generated from some prior distribution that is absolute continuous with respect to Lebesgue measure on $\mathbb{R}$, i.e., $S^{(1)},S^{(2)}\sim p_{s}$, then
@@
Introduction of condition (b).
The condition (b), that the faithfulness condition on $\widetilde{M}$, is to eliminate the cases where though the interventions are applied, it happens that such interventions do not make an impact on the variables intervened. The following example presents such an example.
@@
where $U_{1},U_{2},U_{3}$ are independent standard normal variables, $s\neq 0.5$ is a fixed parameter. We let $e=1$ be the observational environment and $e=2$ be the interventional environment where shifts in mean are applied to the variables $X_{1}$ and $X_{2}$.
@@
A.6 The Complete Statement of Proposition 2